question: Scores on a common final exam are normally distributed with a mean of 66 and a standard deviation of 11. Approximately the top 10% of the students will receive an A. Approximately, what is the minimum exam score to award an A?
I have tried solving this using the standard normal distribution table but have been unable to get the correct answer of 80.

Answer by stanbon(75887)

Scores on a common final exam are normally distributed with a mean of 66 and a standard deviation of 11. Approximately the top 10% of the students will receive an A. Approximately, what is the minimum exam score to award an A?
------
Find the z-value with a left tail of 0.90
------
invNorm(0.90) = 1.2816
----
Find the corresponding score::
x = z*s + u
x = 1.2816*11+66 = 80 when rounded down
----------------
Cheers,
Stan H.
------------------